OpenStudy (anonymous):

if log 2=.301and log 3=.477, what is the approximate value of log 108?

OpenStudy (zolock):

24 = 3 × 8 = 3 × 2³ log 24 = log 3 + 3 log 2 0.477 + 0.903 = 1.38 These are approximate logs to base 10, by the way More accurate value of log10 24 = 1.38021124171

OpenStudy (anonymous):

\[108=2^2\times 3^3\] is a start

OpenStudy (anonymous):



OpenStudy (kc_kennylau):

One formula for you: \(\log a+\log b=\log ab\)

OpenStudy (anonymous):

is the a 2 and the b 3?

OpenStudy (accessdenied):

You should have this in mind... You have approximate values for log 2 and log 3. They ask you to approximate log 108. So, you want to break log 108 apart using logarithm rules so that the only thing left over is log 2 and log 3 in terms of logarithms. Then you can make the approximation happen. example: If you were approximating 144, you could say: 144 = 2*72. Thus, log(144) = log (2*72) = log 2 + log 72. Now you have a log 2 which you know approximates to be .301. Reduce the log(72) further to make more progress...